Q))Two equal chargs are ixed at x=a and x=-a on d axis.anodr pt charge Q is placed at d orign.The change in d electric potential energy of Q,when it is displaced by a small distance x along d x-axis, is approximately proportional to....


(a)x


(b)x2


(c)x3


(d)1/x


Grade:12

1 Answers

Saket Ghanashyam Pande
13 years ago

Initial p.e.=kqQ/a+kqQ/a=2kqQ/a After small displacement p.e= kqQ/a-x + kqQ/x+a = 2kqQa/a2-x2

Net change=2kqQa/a2-x2 - 2kqQ/a = 2kqQx2/a(a2-x2). Since, x is small x2 can be neglected. So, change=2kqQx2/a(a2)

So,change is proportional to x2.

Note: k=1/4π€